OpenStudy (anonymous):

what is the mass, in kilograms, of 14.0 L of gasoline? ( Assume that the density of gasoline is 0.680 g/cm^3) Pleasee help last problem :)))

OpenStudy (unklerhaukus):

you need to know how many cm^3 are in a litre

OpenStudy (anonymous):

1 cm^3 is equal to 1 milliliter, and there are 1000 of those in a liter

OpenStudy (anonymous):

which means there are 14000 cm^3 of the liquid, and then we do proportions

OpenStudy (anonymous):



OpenStudy (anonymous):

well the final answer is 9.52kg

OpenStudy (unklerhaukus):

\[\rho=\frac mV\] \[m=V\rho\] \[V=14.0 [\text L]\times\frac{1000[\text{cm}^3]}{[\text L]}\qquad\qquad\qquad\rho=0.680 \left[\frac{\text g}{\text{cm}^3}\right]\]

OpenStudy (anonymous):

so the mass in grams is 14000*.68 = 9520 grams, and the question asks for kilograms, so divide by 1000 to get 9.52 kilograms

OpenStudy (anonymous):

@mridrik can you rewrite the way you gave me your final answer like write @UnkleRhaukus did?

OpenStudy (anonymous):

I don't know how to type font like that sorry

OpenStudy (anonymous):

just click draw on the bottom of the converstaion box

OpenStudy (anonymous):

right but it would take forever